For the quadratic equation y^2-6y+9=1 find the exact and approximate solutions to three decimal places.
Answer by checkley77(12844)

y^2-6y+9=1
y^2-6y+9-1=0
y^2-6y+8=0
y-4)(y-2)=0
y=4 ans.
y-2=0
y=2 ans.